Pr Nd Ac Th Pa U Np Pu Am Cm BK How many atoms are in 9.61 g of germanium? Select the correct answer below: O 7.97 x 1022
Answer #1

9.60 g Ge = mass / molar mass = 9.60 g / 72.61 g / mole = 0.1322 mole.

1 mole = 6.023 * 10^23

thus

number of atoms = 0.1322 * 6.023 * 10^23 = 7.97 * 10^22 atoms

answer is 7.97 * 10^22 atoms.